y/18=150/126 can you show me how to check it please and thanks
Answer by stanbon(48568)

y/18=150/126
Multiply both sides by 18 to get:
y = 18(150/126)
y = 18(150/(7*18))
Cancel the common 18 in numerator and denominator to get:
y = 150/7
--------------
Checking:
(150/7)/18 = 150/126
Multiply by 18 to get:
150/7 = 150/7
===========
Cheers,
Stan H.